OpenStudy (anonymous):

y = 3x − 5 y = 6x − 8 Explain how you will solve the pair of equations by substitution or elimination. Show all the steps and write the solution

OpenStudy (lynfran):

i would equate the two equation

OpenStudy (lynfran):

3x-5=6x-8

OpenStudy (lynfran):

3x-6x=5-8

OpenStudy (lynfran):

-3x=-3 x=1

OpenStudy (lynfran):

then sub x=1 into the first eq. to find y y=3x-5 y=3(1)-5 y=3-5 y=?

OpenStudy (lynfran):

@hayespurdy

OpenStudy (lynfran):

whats 3-5= to @hayespurdy

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

good so x=1 and y=-2 are the solution ok

OpenStudy (anonymous):

ok thank you about If the two equations are graphed, at what point will the lines representing the two equations intersect?

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

they would intersect at the solution u just found